# A unified approach to hypergeometric class functions

### v1

@@ 7.4 Hermiticity of Sturm-Liouville operators
Then it is easy to see that $\eta_{i}$ are real, and if $\eta_{1}\neq\eta_{2}$, then $(f_{i}|f_{j})=0$. Therefore, eigenfunctions of Hermitian operators can be arranged into orthogonal families. If an operator is Hermitian and possesses an orthogonal basis of eigenvectors, then it is is essentially self-adjoint on linear combinations of its eigenvectors.
@@ 7.5 Selecting endpoints for Sturm-Liouville operators
7.6 Orthogonality of hypergeometric polynomials
In some cases hypergeometric polynomials can be interpreted as an orthogonal basis of a certain weighted Hilbert space. In this case they are often called (very) classical orthogonal polynomials.
@@ 7.6 Orthogonality of hypergeometric polynomials
${\cal A}$ $:=\sigma(x)\partial_{x}^{2}+\tau(x)\partial_{x}=\rho(x)^{-1}\partial_{x}\rho(x)\sigma(x)\partial_{x}.$ (7.29)
is Hermitian on the space of polynomials in the sense of the Hilbert space $L^{2}(]a,b[,\rho)$. Therefore, polynomial eigenfunctions of ${\cal A}$ are pairwise orthogonal (at least if the corresponding eigenvalues are distinct, which is generically the case, see (7.8)). The following theorem says more: we compute the square norm of the polynomials obtained from the Rodriguez formula.
@@ 7.7 Review of types of hypergeometric class polynomials
In the remaining part of the section we review various types of hypergeometric polynomials. We will discuss only the properties that follow directly from the general theory described in previous subsections. Here is the list of items that we will cover:
@@
Let us mention that hypergeometric class polynomials possess various useful features that will not be listed in the following subsections. For instance, they typically have additional recurrence relations and generating functions. We list only those that follow directly from the general theory described above.
@@ 7.8 Jacobi polynomials
    If $\alpha+\beta\not\in\{-2n,\dots,-n-1\}$, then $\deg P_{n}^{\alpha,\beta}=n$. It is then up to a coefficient the unique eigensolution of the operator
    ${\cal A}:=(1-x^{2})\partial_{x}^{2}+(\beta-\alpha-(\alpha+\beta+2)x)\partial_{x},$ (7.38)
@@
If $\alpha,\beta>-1$, then ${\cal A}$ is self-adjoint on the space of polynomials in the sense of
@@
The Jacobi polynomials are its eigenfunctions and form an orthogonal basis with the normalization
$\int_{-1}^{1}(P_{n}^{\alpha,\beta}(x))^{2}(1-x)^{\alpha}(1+x)^{\beta}{\rm d}x=\frac{\Gamma(1+\alpha+n)\Gamma(1+\beta+n)2^{\alpha+\beta+1}}{(1+2n+\alpha+\beta)n!\Gamma(1+\alpha+\beta+n)}.$ (7.42)
@@ 7.9 Laguerre polynomials
${\cal A}:=x\partial_{x}^{2}+(\alpha+1-x)\partial_{x},$ (7.44)
@@
If $\alpha>-1$, then ${\cal A}$ is essentially self-adjoint on the space of polynomials in the sense of $L^{2}([0,\infty[,{\rm e}^{-x}x^{\alpha})$. The Laguerre polynomials are its eigenfunctions and form an orthonormal basis with the normalization
@@ 7.10 Bessel polynomials
${\cal A}:=z^{2}\partial_{z}^{2}+(-1+(2+\theta)z)\partial_{z},$ (7.52)
@@ 7.11 Hermite polynomials
${\cal A}:=\partial_{x}^{2}-2x\partial_{x},$ (7.57)
@@
The operator ${\cal A}$ is essentially self-adjoint on the space of polynomials in the sense of $L^{2}({\mathbb{R}},{\rm e}^{-x^{2}})$. The Hermite polynomials are its eigenfunctions and form an orthogonal basis with the normalization
@@ 8 Degenerate case
If two indices of a regular-singular (also called Fuchsian) point of a differential equation differ by an integer, then the usual Frobenius method [WW] produces in general only one solution, up to a coefficient. We call this case degenerate. In this section we will be discuss the degenerate case of hypergeometric class equations. Without limiting the generality, the regular-singular point will be 00.
8.1 United hypergeometric function in the degenerate case
@@ 8.3 Generating functions
Let us now fix $\sigma(z)=\frac{\sigma^{\prime\prime}}{2}z^{2}+z$, as in (8.1). Fix also $\delta,\theta$. Let $\rho$ and $\tilde{\rho}$ solve the equations
$-\frac{\sigma^{\prime\prime}}{2}\delta=\rho\tilde{\rho},\qquad\rho+\tilde{\rho}=\theta.$ (8.8)
Consider two families of hypergeometric class operators
$\Big{(}\frac{\sigma^{\prime\prime}}{2}z^{2}+z\Big{)}\partial_{z}^{2}+\Big{(}m+1+\Big{(}\frac{\sigma^{\prime\prime}}{2}(1+m)-\theta\Big{)}z\Big{)}\partial_{z}-m\rho-\delta,$ (8.9)
$\Big{(}\frac{\sigma^{\prime\prime}}{2}z^{2}+z\Big{)}\partial_{z}^{2}+\Big{(}m+1+\Big{(}\frac{\sigma^{\prime\prime}}{2}(1+m)-\theta\Big{)}z\Big{)}\partial_{z}-m\tilde{\rho}-\delta.$ (8.10)
In terms of our standard parameters, these operators can be written as
${\mathcal{C}}(\sigma,\kappa_{m})+\omega_{m},$ (8.11)
${\mathcal{C}}(\sigma,\kappa_{m})+\tilde{\omega}_{m},$ (8.12)
@@
$\kappa_{m}(z)$ $:=m+\Big{(}\frac{\sigma^{\prime\prime}}{2}(m-1)-\theta\Big{)}z,$ (8.13)
$\omega_{m}$ $:=\frac{\theta}{2}-\delta-m\rho-\frac{\sigma^{\prime\prime}}{4}(m-1),$ (8.14)
$\tilde{\omega}_{m}$ $:=\frac{\theta}{2}-\delta-m\tilde{\rho}-\frac{\sigma^{\prime\prime}}{4}(m-1).$ (8.15)
@@
We are going to construct solutions to the equations given by (8.9) and (8.10). In order to present them in a unified way, it is convenient to choose $a,b,\mu,\nu$ satisfying
$\rho=\mu b,\quad\tilde{\rho}=\nu a,\quad\delta=ab.$ (8.17)
By $[(\alpha,\beta)^{+}]$ we will denote the counterclockwise loop around $\alpha,\beta\in{\mathbb{C}}$. We define
@@
    The functions $\Psi_{m}(z)$ and $\tilde{\Psi}_{m}(z)$ are proportional to hypergeometric functions:
@@
Proof.  0. follows immediately from the definition and the Laurent expansion.
We can rewrite (8.20) as
$(1+\mu u)^{-\frac{a}{\mu}}(1+\nu v)^{-\frac{b}{\nu}}=\sum_{m\in{\mathbb{Z}}}u^{m}\Psi_{m}(uv).$ (8.29)
By (8.21) we have
@@
This implies 1. 2. is immediate.
Let us show 3.. We have the identity
$\begin{array}[]{rl}0&=\Big{(}\partial_{u}((1+\nu v)\partial_{v}+b)+\partial_{v}((1+\mu u)\partial_{u}+a)\\[8.53581pt] &-((1+\mu u)\partial_{u}+u)((1+\nu v)\partial_{v}+b)\Big{)}(1+\mu u)^{-\frac{a}{\mu}}(1+\nu v)^{-\frac{b}{\nu}}\\[8.53581pt] &=\left(\partial_{u}\partial_{v}-\mu\nu uv\partial_{u}\partial_{v}-\mu bu\partial_{u}-a\nu v\partial_{v}-ab\right)(1+\mu u)^{-\frac{a}{\mu}}(1+\nu v)^{-\frac{b}{\nu}}.\end{array}$ (8.31)
Besides, if $z=uv$, then
$u\partial_{u}u^{m}\Psi_{m}(uv)=u^{m}(m+z\partial_{z})\Psi_{m}(z),$
$v\partial_{v}u^{m}\Psi_{m}(uv)=u^{m}z\partial_{z}\Psi_{m}(z),$
$\partial_{v}\partial_{u}u^{m}\Psi_{m}(uv)=u^{m}\left((m+1)\partial_{z}+z\partial_{z}^{2}\right)\Psi_{m}(z),$
Hence
@@
Thus
$0=\sum_{m\in{\mathbb{Z}}}u^{m}\left(z(1-\mu\nu z)\partial_{z}^{2}+(m+1-(\mu\nu+(m\mu+a)\nu+b\mu)z)\partial_{z}-(m\mu+a)b\right)\Psi_{m}(z),$
which implies (8.25). $\Box$
@@ 8.6 The ${}_{0}F_{1}$ function
Acknowledgement. This work was supported by National Science Center (Poland) under the grant UMO-2019/35/B/ST1/01651. I thank Rupert Frank for an inspiring remark. I am also grateful to Adam Kaliński for useful comments on the manuscript.

### v2

@@ 7.4 Hermiticity of Sturm-Liouville operators
Then it is easy to see that $\eta_{i}$ are real, and if $\eta_{1}\neq\eta_{2}$, then $(f_{i}|f_{j})=0$. Therefore, eigenfunctions of Hermitian operators can be arranged into orthogonal families. If an operator is Hermitian and possesses an orthogonal basis of eigenvectors, then it is is essentially self-adjoint on (finite) linear combinations of its eigenvectors.
@@ 7.5 Selecting endpoints for Sturm-Liouville operators
7.6 Orthogonality of hypergeometric class polynomials
In some cases hypergeometric class polynomials can be interpreted as an orthogonal basis of a certain weighted Hilbert space. They are then often called (very) classical orthogonal polynomials.
@@ 7.6 Orthogonality of hypergeometric class polynomials
${\cal H}(\sigma,\kappa)$ $:=\sigma(x)\partial_{x}^{2}+\big{(}\kappa(x)+\sigma^{\prime}(x)\big{)}\partial_{x}+\frac{\kappa^{\prime}}{2}=\rho(x)^{-1}\partial_{x}\rho(x)\sigma(x)\partial_{x}+\frac{\kappa^{\prime}}{2}.$ (7.29)
is Hermitian on the space of polynomials in the sense of the Hilbert space $L^{2}(]a,b[,\rho)$. Therefore, polynomial eigenfunctions of ${\cal H}(\sigma,\kappa)$ are pairwise orthogonal (at least if the corresponding eigenvalues are distinct, which is generically the case, see (7.8)). The following theorem says more: we compute the square norm of the polynomials obtained from the Rodriguez formula.
@@ 7.7 Review of types of hypergeometric class polynomials
In the remaining part of the section we review various types of hypergeometric class polynomials. We will discuss only the properties that follow directly from the general theory described in previous subsections. Here is the list of items that we will cover:
@@
Hypergeometric class polynomials possess various useful features that will not be listed in the following subsections. For instance, they typically have additional recurrence relations and generating functions. We list only those that follow directly from the general theory described above.
@@ 7.8 Jacobi polynomials
    If $\alpha+\beta\not\in\{-2n,\dots,-n-1\}$, then $\deg P_{n}^{\alpha,\beta}=n$. It is then up to a coefficient the unique eigenfunction of the operator
    ${\cal H}(\sigma,\kappa)=(1-x^{2})\partial_{x}^{2}+(\beta-\alpha-(\alpha+\beta+2)x)\partial_{x}+\frac{\beta-\alpha}{2}$ (7.38)
@@
If $\alpha,\beta>-1$, then ${\cal H}(\sigma,\kappa)$ is self-adjoint on the space of polynomials in the sense of
@@
Jacobi polynomials are its eigenfunctions and form an orthogonal basis with the normalization
$\int_{-1}^{1}P_{n}^{\alpha,\beta}(x)^{2}(1-x)^{\alpha}(1+x)^{\beta}{\rm d}x=\frac{\Gamma(1+\alpha+n)\Gamma(1+\beta+n)2^{\alpha+\beta+1}}{(1+2n+\alpha+\beta)n!\Gamma(1+\alpha+\beta+n)}.$ (7.42)
@@ 7.9 Laguerre polynomials
${\cal H}(\sigma,\kappa)=x\partial_{x}^{2}+(\alpha+1-x)\partial_{x}-\frac{1}{2}$ (7.44)
@@
If $\alpha>-1$, then ${\cal H}(\sigma,\kappa)$ is essentially self-adjoint on the space of polynomials in the sense of $L^{2}([0,\infty[,{\rm e}^{-x}x^{\alpha})$. Laguerre polynomials are its eigenfunctions and form an orthonormal basis with the normalization
@@ 7.10 Bessel polynomials
${\cal H}(\sigma,\kappa)=z^{2}\partial_{z}^{2}+(-1+(2+\theta)z)\partial_{z}+\frac{\theta}{2}$ (7.52)
@@ 7.11 Hermite polynomials
${\cal H}(\sigma,\kappa)=\partial_{x}^{2}-2x\partial_{x}-1$ (7.57)
@@
The operator ${\cal H}(\sigma,\kappa)$ is essentially self-adjoint on the space of polynomials in the sense of $L^{2}({\mathbb{R}},{\rm e}^{-x^{2}})$. Hermite polynomials are its eigenfunctions and form an orthogonal basis with the normalization
@@ 8 Degenerate case
If two indices of a regular-singular (also called Fuchsian) point of a differential equation differ by an integer, then the usual Frobenius method [WW] produces in general, up to a coefficient, only one solution. We call this case degenerate. In this section we will be discuss the degenerate case of hypergeometric class equations. Without limiting the generality, the regular-singular point will be 00.
8.1 Unified hypergeometric function in the degenerate case
@@ 8.3 Generating functions
Let us now fix $\sigma(z)=\frac{\sigma^{\prime\prime}}{2}z^{2}+z$, as in (8.1). Let $a,b,\mu,\nu\in{\mathbb{C}}$. Consider two families of hypergeometric class operators
$\Big{(}\frac{\sigma^{\prime\prime}}{2}z^{2}+z\Big{)}\partial_{z}^{2}+\Big{(}m+1+\Big{(}\frac{\sigma^{\prime\prime}}{2}(1+m)-\mu b-\nu a\Big{)}z\Big{)}\partial_{z}-m\mu b-ab,$ (8.8)
$\Big{(}\frac{\sigma^{\prime\prime}}{2}z^{2}+z\Big{)}\partial_{z}^{2}+\Big{(}m+1+\Big{(}\frac{\sigma^{\prime\prime}}{2}(1+m)-\mu b-\nu a\Big{)}z\Big{)}\partial_{z}-m\nu a-ab.$ (8.9)
(Note that the parameters we introduced are redundant—three would suffice instead of four. However they help us to describe the degenerate case in a symmetric and unified way).
In terms of our standard parameters, the operators (8.8) and (8.9) can be written as
${{\cal H}}(\sigma,\kappa_{m})+\omega_{m},$ (8.10)
${{\cal H}}(\sigma,\kappa_{m})+\tilde{\omega}_{m},$ (8.11)
@@
$\kappa_{m}(z)$ $:=m+\Big{(}\frac{\sigma^{\prime\prime}}{2}(m-1)-\mu b-\nu a\Big{)}z,$ (8.12)
$\omega_{m}$ $:=\frac{1}{2}(\mu b+\nu a)-ab-m\mu b-\frac{\sigma^{\prime\prime}}{4}(m-1),$ (8.13)
$\tilde{\omega}_{m}$ $:=\frac{1}{2}(\mu b+\nu a)-ab-m\nu a-\frac{\sigma^{\prime\prime}}{4}(m-1).$ (8.14)
@@
By $[(\alpha,\beta)^{+}]$ we will denote a counterclockwise loop around $\alpha,\beta\in{\mathbb{C}}$. We define
@@
    The functions $\Psi_{m}(z)$ and $\tilde{\Psi}_{m}(z)$ are proportional to the Olver normalized unified hypergeometric function:
@@
Proof.  1. follows immediately from the definitions (8.16), (8.17) and the Laurent expansion.
To show 2. we can rewrite (8.18) and (8.19) as
$(1+\mu u)^{-\frac{a}{\mu}}(1+\nu v)^{-\frac{b}{\nu}}$ $=\sum_{m\in{\mathbb{Z}}}u^{m}\Psi_{m}(uv),$ (8.26)
@@
3. follows by setting $v=0$ in (8.26) and $u=0$ in (8.27).
Let us show 4. We have the identity
00 0 $=\left(\partial_{u}\partial_{v}-\mu\nu uv\partial_{u}\partial_{v}-\mu bu\partial_{u}-a\nu v\partial_{v}-ab\right)(1+\mu u)^{-\frac{a}{\mu}}(1+\nu v)^{-\frac{b}{\nu}}.$ (8.28)
Besides,
@@
Thus (8.28) together with (8.26) can be rewritten as
$0=\sum_{m\in{\mathbb{Z}}}u^{m}\left(z(1-\mu\nu z)\partial_{z}^{2}+(m+1-(\mu\nu+(m\mu+a)\nu+b\mu)z)\partial_{z}-(m\mu+a)b\right)\Psi_{m}(z),$ (8.29)
which implies (8.22).
5. follows from 3. and 4. $\Box$
@@ 8.6 The ${}_{0}F_{1}$ function
Acknowledgement. This work was supported by National Science Center (Poland) under the grant UMO-2019/35/B/ST1/01651. I thank Rupert Frank for an inspiring remark. I am also grateful to Adam Kaliński for useful comments on the manuscript. I have also greatly profited from many helpful remarks of the referee of the previous version of the manuscript.
Appendix A Function ${}_{2}F_{0}$
The function ${}_{2}F_{0}$ seems to be rarely discussed in the literature. For convenience of the reader in the following theorem we state and prove basic facts about this function.
Theorem A.1
For $w\in{\mathbb{C}}\backslash[0,\infty[$, there exists the limit
$F(a,b;-;w):=\lim_{c\to+\infty}F(a,b;c;cw).$ (A.1)
It defines a function analytically depending on $a,b\in{\mathbb{C}}$ and $w\in{\mathbb{C}}\backslash[0,\infty[$.
We have the following asymptotic expansion:
$F(a,b;-;w)\sim\sum_{n=0}^{\infty}\frac{(a)_{n}(b)_{n}}{n!}w^{n}.$ (A.2)
More precisely, for any $\epsilon>0$, $n$, there exists $c_{n}$ such that
$\Big{|}F(a,b;-;w)-\sum_{j=0}^{n}\frac{(a)_{j}(b)_{j}}{j!}w^{j}\Big{|}\leq c_{n}|w|^{n+1},\quad|\arg w|\geq\epsilon,\quad|w|<1.$ (A.3)
Moreover, for $\mathrm{Re}(a)>0$ we have an integral representation
$F(a,b;-;w)=\frac{1}{\Gamma(a)}\int_{0}^{\infty}{\rm e}^{-t}t^{a-1}(1-wt)^{-b}{\rm d}t.$ (A.4)
Proof.  Assume first that $\mathrm{Re}(a)>0$. For $\mathrm{Re}(c-a)>0$ and $w\in{\mathbb{C}}\backslash[c^{-1},+\infty[$, inserting $t=cs^{-1}$ and $z=cw$ into (6.1), we obtain
$F(a,b;c;cw)=\frac{\Gamma(c)c^{-a}}{\Gamma(a)\Gamma(c-a)}\int_{0}^{c}s^{a-1}(1-c^{-1}s)^{c-a-1}(1-ws)^{-b}{\rm d}s.$ (A.5)
Using $\lim_{c\to\infty}\frac{\Gamma(c)c^{-a}}{\Gamma(c-a)}=1$ and the Lebesgue Dominated Convergence Theorem we see that (A.5) converges to the right hand side of (A.4). This proves that $F(a,b;-;w)$ is well defined for $\mathrm{Re}(a)>0$.
Now let $a$ be arbitrary. We have
$\partial_{z}^{n}F(a,b;c;z)=\frac{(a)_{n}(b)_{n}}{(c)_{n}}F(a+n,b+n;c+n;z).$ (A.6)
Using Taylor’s formula with a remainder
$f(z)=\sum_{j=0}^{n-1}\frac{f^{(j)}(0)z^{j}}{j!}+z^{n}\int_{0}^{1}\frac{f^{(n)}(sz)n(1-s)^{n-1}}{n!}{\rm d}s,$
we obtain
$F(a,b;c;z)$ $=\sum_{j=0}^{n-1}\frac{(b)_{j}(a)_{j}w^{j}}{j!}$ (A.7)
$+\frac{w^{n}(b)_{n}(a)_{n}}{(n-1)!}\int_{0}^{1}(1-s)^{n-1}{\rm d}sF(a+n,b+n;c;ws).$ (A.8)
Now, choose $n$ such that $\mathrm{Re}(a+n)>0$. Then we can apply what we proved before to show the convergence
$\lim_{c\to+\infty}F(a,b;c;cz)$ $=\sum_{j=0}^{n-1}\frac{(b)_{j}(a)_{j}w^{j}}{j!}$ (A.9)
$+\frac{w^{n}(b)_{n}(a)_{n}}{(n-1)!}\int_{0}^{1}(1-s)^{n-1}{\rm d}sF(a+n,b+n;-;ws).$ (A.10)
Clearly, (A.10) is $O(w^{n})$, and $n$ can be made arbitrarily large. $\Box$
If $a$ or $b$ is a negative integer, then the series (A.2) terminates and functions ${}_{2}F_{0}(a,b;-;\cdot)$ essentially coincide with Bessel polynomials, see Subsect. 7.10. Otherwise, ${}_{2}F_{0}(a,b;-;\cdot)$ have a logarithmic branch point at 00.
Appendix B Riemann class
We hope that we convinced our reader that the hypergeometric class studied in this paper is a natural family of equations. In this appendix, following [DIL], we will describe a somewhat wider family, which appears in the literature in many sources, e.g. [SL, DIL, EEKS]. According to the terminology used in [DIL], directly inspired by and partly borrowed from the monograph of Slavyanov-Lay [SL], this family is called Riemann class.
Consider an equation given by the operator
$\partial_{z}^{2}+b(z)\partial_{z}+c(z),$ (B.1)
where $p(z)$, $q(z)$ are rational functions. Let $z_{0}\in{\mathbb{C}}\cup\{\infty\}$ be a singular point of (B.1).
Recall that $z_{0}\in{\mathbb{C}}$ is called regular-singular or Fuchsian if $b(z)=\frac{p(z)}{z-z_{0}}$ and $c(z)=\frac{q(z)}{(z-z_{0})^{2}}$ with $p,q$ regular at $z_{0}$. The equation
$\lambda(\lambda-1)+p(z_{0})\lambda+q(z_{0})=0$ (B.2)
is called the indicial equation of $z_{0}$ and its roots are called indices of $z_{0}$.
We say that $\infty$ is regular-singular or Fuchsian if $b(z)=\frac{p(z)}{z}$ and $c(z)=\frac{q(z)}{z^{2}}$ with $p,q$ regular at $\infty$.
$\lambda(\lambda+1)-p(\infty)\lambda+q(\infty)=0$ (B.3)
is called the indicial equation of $\infty$ and its roots are called indices of $\infty$.
It is easy to see that every equation having no more than $3$ singular points in ${}{\mathbb{C}}\cup\{\infty\}$, all of them Fuchsian and at most $2$ finite, is given by an operator of the form
$\partial_{z}^{2}+\Big{(}\frac{a_{1}}{z-z_{1}}+\frac{a_{2}}{z-z_{2}}\Big{)}\partial_{z}+\frac{b_{1}}{z-z_{1}}+\frac{b_{2}}{z-z_{2}}+\frac{c_{1}}{(z-z_{1})^{2}}+\frac{c_{2}}{(z-z_{2})^{2}}\quad\text{ with }\quad b_{1}+b_{2}=0,$ (B.4)
where $z_{1},z_{2}$ are distinct points in ${\mathbb{C}}$. Following [DIL], the family of equations (B.4) is called the Riemann type. (Another name, introduced in [SL], is $M_{2}$-type).
Each finite singularity has at least one index equal 00 if and only if $c_{1}=c_{2}=0.$ Such equations are given by operators
$\partial_{z}^{2}+\Big{(}\frac{a_{1}}{z-z_{1}}+\frac{a_{2}}{z-z_{2}}\Big{)}\partial_{z}+\frac{b_{1}}{z-z_{1}}+\frac{b_{2}}{z-z_{2}}\quad\text{ with }\quad\sum_{j=1}^{n}b_{j}=0.$ (B.5)
Following [DIL], the family of equations given by (B.5) will be called the grounded Riemann type. It is easy to see that by gauging with power functions we can always transform a Riemann type equation into a grounded Riemann type equation.
We say that a differential equation belongs to the Riemann class (or the $M_{2}$-class) if it is given by
$\partial_{z}^{2}+\frac{\tau(z)}{\sigma(z)}\partial_{z}+\frac{\xi(z)}{\sigma(z)^{2}},$ (B.6)
where $\sigma,\tau,\xi$ are polynomials satisfying
$\sigma\neq 0,\quad\deg\sigma\leq 2,\quad\deg\tau\leq 1,\quad\deg\xi\leq 2.$ (B.7)
Thus the Riemann class comprises the Riemann type together with all its confluent cases.
We say that a differential equation belongs to the grounded Riemann class if it is given by
$\partial_{z}^{2}+\frac{\tau(z)}{\sigma(z)}\partial_{z}+\frac{\eta(z)}{\sigma(z)},$ (B.8)
where $\sigma,\tau,\eta$ are polynomials satisfying
$\sigma\neq 0,\quad\deg\sigma\leq 2,\quad\deg\tau\leq 1,\quad\deg\eta=0.$ (B.9)
Thus the grounded Riemann class besides the grounded Riemann includes all its confluent cases.
The following proposition is proven in [DIL]:
Proposition B.1
  1. 1.
    The Riemann type is contained in the Riemann class. An equation of the Riemann class is of the Riemann type iff $\sigma$ possesses $2$ distinct roots.
  2. 2.
    The grounded Riemann type is the intersection of the grounded Riemann class and Riemann type.
In this paper instead of the grounded Riemann class we use the name hypergeometric class. We use this name mostly for brevity, besides the word “grounded” may sound bizarre to some readers. Furthermore, we prefer to multiply these equations by $\sigma(z)$, so that we consider operators of the form
$\sigma(z)\partial_{z}^{2}+\tau(z)\partial_{z}+\eta$ (B.10)
with $\sigma,\tau,\nu$ satisfying (B.9).
One can ask whether the unified theory presented in this paper can be extended to operators of the form
$\sigma(z)\partial_{z}^{2}+\tau(z)\partial_{z}+\frac{\xi(z)}{\sigma(z)},$ (B.11)
where $\sigma,\tau,\xi$ satisfy (B.7). In other words whether the unified theory presented in this paper can be extended to the full Riemann class. We do not know an answer to this question. However, let us remark that we do not gain much by considering the full Riemann class instead of the grounded Riemann class.
By a division by a constant, transformation $z\mapsto az+b$ and gauging with a power and exponential all operators from the Riemann type can be always reduced to an ${}_{2}F_{1}$ operator. More generally, all operators from the Riemann class can be reduced to one of normal forms listed in [SL] or in Subsect. 3.3 of [DIL]:
the ${}_{2}F_{1}$ operator $z(1-z)\partial_{z}^{2}+\big{(}c-(a+b+1)z\big{)}\partial_{z}-ab$;
the ${}_{2}F_{0}$ operator $z^{2}\partial_{z}^{2}+\big{(}-1+(a+b+1)z\big{)}\partial_{z}+ab$;
the ${}_{1}F_{1}$ operator $z\partial_{z}^{2}+(c-z)\partial_{z}-a$;
the ${}_{0}F_{1}$ operator $z\partial_{z}^{2}+c\partial_{z}-1$;
the Hermite operator $\partial_{z}^{2}-2z\partial_{z}-2a$;
the Airy operator $\partial_{z}^{2}+z$;
the Euler operator I $z^{2}\partial_{z}^{2}+cz\partial_{z}$;
the Euler operator II $z\partial_{z}^{2}+c\partial_{z}$;
the 1d Helmholtz operator $\partial_{z}^{2}+1$;
the 1d Laplace operator $\partial_{z}^{2}$.
In this list the first five are the normal forms listed in the introduction and described in Section 6. The last four are trivial operators that can be solved in by elementary methods. They are also special cases of the grounded Riemann class.
The Airy equation is the only type within the Riemann class which does not belong to the grounded Riemann class and therefore is left out from the unified theory presented in this paper.